# Existence, uniqueness, and time-asymptotics of regular solutions in multidimensional thermoelasticity on domains with boundary

### v1

@@ 5 Time-asymptotics of solutions
On the other hand, the rotation-free part of the displacement and the temperature are coupled in the same system as 1.1. The potential part and $\theta$ behave in the same way as the heated string in [5]. Namely, the curl-free part is influenced by the propagation of heat. It means that the whole energy, interlocked with the displacement, is dissipated into heat as $t\to\infty$.
@@ 5.2 Displacement and temperature at infinity
It occurs that $\boldsymbol{\gamma}$ and $\boldsymbol{\chi}$ behave differently when $t\to\infty$. We start by noticing that the non-divergence part of $\mathbf{u}$ (i.e., $\boldsymbol{\gamma}$) oscillates unless it is zero (cf.[3]).
@@
Let us now study the asymptotics of the potential part of $\mathbf{u}$ (i.e., $\boldsymbol{\chi}$) and the temperature. Our considerations are similar to those in [5, 6]. Nevertheless, certain details need to be converted. We begin with the following proposition regarding the stability of solutions.

### v2

@@ 5 Time-asymptotics of solutions
On the other hand, the rotation-free part of the displacement and the temperature are coupled in the same system as 1.1. The potential part and $\theta$ behave in the same way as the heated string in BC2 . Namely, the curl-free part is influenced by the propagation of heat. It means that the whole energy, interlocked with the displacement, is dissipated into heat as $t\to\infty$.
@@ 5.2 Displacement and temperature at infinity
It occurs that $\boldsymbol{\gamma}$ and $\boldsymbol{\chi}$ behave differently when $t\to\infty$. We start by noticing that the non-divergence part of $\mathbf{u}$ (i.e., $\boldsymbol{\gamma}$) oscillates unless it is zero (cf.Bies ).
@@
Let us now study the asymptotics of the potential part of $\mathbf{u}$ (i.e., $\boldsymbol{\chi}$) and the temperature. Our considerations are similar to those in BC2 ; Biest . Nevertheless, certain details need to be converted. We begin with the following proposition regarding the stability of solutions.